The Human Toxome Project (HTP) is working to understand the scope of industrial pollution in thehuman body. Industrial chemicals may enter the body through pollution or as ingredients inconsumer products. In October 2008, the scientists at HTP tested cord blood samples for 20newborn infants in the United States. The cord blood of the "In utero/newborn" group wastested for 140 industrial compounds, pollutants, and other chemicals, including chemicals linkedto brain and nervous system toxicity, immune system toxicity, and reproductive toxicity, andfertility problems. There are health concerns about the effects of some chemicals on the brainand nervous system.  Here is the data:

79, 145,147,160, 116, 100, 159, 151, 156, 126, 137, 83, 156, 94, 121, 144, 123, 114, 139, 99

At the 5% level of significance, do newborn infants have less than 140 compounds in their cord blood?

This problem is a one-mean t-test because:

Question 1 options:

There is a population standard deviation that is known.

The sample standard deviation is known but no population standard deviation is known.

There is a level of significance that is known.

Question 2 (1 point)

What is the population mean?

Question 2 options:

140 compounds

150 compounds

127 compounds

Question 3 (1 point)

What is the sample mean?

Question 3 options:

150 compounds

127 compounds

140 compounds

Question 4 (1 point)

What is the sample standard deviation?

Question 4 options:

26 compounds

20 compounds

127 compounds

Question 5 (1 point)

What is the level of significance?

Question 5 options:

0.05

0.01

0.50

Question 6 (1 point)

What is the sample size?

Question 6 options:

10

15

20

Question 7 (1 point)

Choose the correct null hypothesis (Ho):

Question 7 options:

Ho: µ ≥ 140 compounds

Ho: µ ≤ 140 compounds

Ho: µ = 140 compounds

Question 8 (1 point)

Choose the correct alternative hypothesis (Ha):

Question 8 options:

Ha: µ < 140 compounds

Ha: µ > 140 compounds

Ha: µ ≠ 140 compounds

Question 9 (1 point)

What is the T- test statistics?

Question 9 options:

0

2.16

-2.16

Question 10 (1 point)

What is the p-value?

Question 10 options:

.022

.22

.0022

Question 11 (1 point)

What is the conclusion when you compare the p-value and the level of significance (alpha)?

Question 11 options:

Reject the null since the p-value is equal to the level of significance.

Reject the null since the p-value is less than the level of significance.

Fail to Reject the null since the p-value is greater than the level of significance.

Question 12 (1 point)

What is the conclusion based on the original claim?

Question 12 options:

At the 5% level of significance, newborn infants do not have less than 140 compounds in their cord blood.

At the 5% level of significance, newborn infants have less than 140 compounds in their cord blood.
Answer #1

Question 1

This problem is a one-mean t-test because:

The sample standard deviation is known but no population standard deviation is known

Question 2

What is the population mean?

140 compounds

Question 3

What is the sample mean

Given,

\overline{x}=\frac{\Sigma x}{n}

n : sample size = 20

x:Number of compunds
79
145
147
160
116
100
159
151
156
126
137
83
156
94
121
144
123
114
139
99
Total \Sigma x=2549
Mean: \overline{x} =2549/20 127.45

Ans :

127 compounds

Question 4

What is the sample standard deviation?

S Σ(x – )2 TV η -1 = 1

x:Number of compunds (x-\overline{x}) (x-\overline{x})2
79 -48.45 2347.4025
145 17.55 308.0025
147 19.55 382.2025
160 32.55 1059.5025
116 -11.45 131.1025
100 -27.45 753.5025
159 31.55 995.4025
151 23.55 554.6025
156 28.55 815.1025
126 -1.45 2.1025
137 9.55 91.2025
83 -44.45 1975.8025
156 28.55 815.1025
94 -33.45 1118.9025
121 -6.45 41.6025
144 16.55 273.9025
123 -4.45 19.8025
114 -13.45 180.9025
139 11.55 133.4025
99 -28.45 809.4025
Total 2549 Σ(x – ΤΣ=12808.95
Mean 127.45

Σ(A - Z)2 η 1 12808.95 -0 90 1 O-V = 25.9645

Ans : 26

What is the level of significance?

Ans: 0.05

Question 6 (1 point)

What is the sample size?

Ans : 20

Question 7

Choose the correct null hypothesis (Ho):

Ho: \mu = 140 compounds

Question 8

Choose the correct alternative hypothesis (Ha):

Ha : \mu < 140 compounds

Question 9 (1 point)

What is the T- test statistics

T-HO (127.45 – 140) Test Statistic : Istat = s/n = 25.9645/20 -12.55 - 5.8058 25.9645/20 = 5.8058 = -2.1616

Ans:

-2.16

Question 10 (1 point)

What is the p-value?

for left tailed test:

P - Value = P(t <tstat) = P(t < -2.1616)

for 19 degrees of freedom,  P(t<-2.1616)=0.0218

p-value = 0.0218

p-value = 0.022

Question 11 (1 point)

What is the conclusion when you compare the p-value and the level of significance (alpha)?

Reject the null since the p-value is less than the level of significance.

Question 12 (1 point)

What is the conclusion based on the original claim?

At the 5% level of significance, newborn infants have less than 140 compounds in their cord blood.